An 18-foot ladder is placed 6 feet from the wall. Approximately how high does the ladder reach?
Answer by nerdybill(5411)

An 18-foot ladder is placed 6 feet from the wall. Approximately how high does the ladder reach?
.
Draw a diagram of the problem, it'll help you "see" the solution easier.
.
Basically, you have a right-triangle so you can apply Pythagorean theorem:
.
Let h = height the ladder reaches
then
h^2 + 6^2 = 18^2
h^2 = 18^2 - 6^2
h^2 = 324 - 36
h^2 = 288
h = sqrt(288)
h = 16.97 feet